# More Experts Than Galaxies: Conditionally-Overlapping Experts with Biologically-Inspired Fixed Routing

### v2

@@ Abstract
The evolution of biological neural systems has led to both modularity and sparse coding, which enables efficiency in energy usage, and robustness across the diversity of tasks in the lifespan. In contrast, standard neural networks rely on dense, non-specialized architectures, where all model parameters are simultaneously updated to learn multiple tasks, leading to representation interference. Current sparse neural network approaches aim to alleviate this issue, but are often hindered by limitations such as 1) trainable gating functions that cause representation collapse; 2) non-overlapping experts that result in redundant computation and slow learning; and 3) reliance on explicit input or task IDs that impose significant constraints on flexibility and scalability. In this paper we propose Conditionally Overlapping Mixture of ExperTs (COMET), a general deep learning method that addresses these challenges by inducing a modular, sparse architecture with an exponential number of overlapping experts. COMET replaces the trainable gating function used in Sparse Mixture of Experts with a fixed, biologically inspired random projection applied to individual input representations. This design causes the degree of expert overlap to depend on input similarity, so that similar inputs tend to share more parameters. This facilitates positive knowledge transfer, resulting in faster learning and improved generalization. We demonstrate the effectiveness of COMET on a range of tasks, including image classification, language modeling, and regression, using several popular deep learning architectures.
@@ 1 INTRODUCTION
In recent years, there has been a trend towards developing increasingly larger models (OpenAI, 2023a; b; Fedus et al., 2022; Shuster et al., 2022; Chowdhery et al., 2022), driven by the understanding that a neural network’s learning capacity depends on its number of parameters (Shazeer et al., 2017). This approach has yielded impressive results in various fields, including computer vision (Dosovitskiy et al., 2021; Kirillov et al., 2023) and language modeling (OpenAI, 2023b; Chowdhery et al., 2022). However, with such large size come difficulties, including increased training costs, decreased interpretability, and growing requirements for large amounts of memory and storage.
@@
One approach to mitigating some of these challenges is sparsity, where a subset of the model’s parameters is selectively utilized in the computational graph. This concept of sparsity has been widely explored in machine learning (Jacobs et al., 1991; Jordan & Jacobs, 1993; LeCun et al., 1989; Zhou et al., 2019; Hoefler et al., 2021; Shazeer et al., 2017; Bair et al., 2024). Researchers have observed significant benefits of sparsity, including reduced inference costs (Han et al., 2016; Shazeer et al., 2017), improved generalization capabilities (LeCun et al., 1989; Jacobs & Burkholz, 2024; Frankle & Carbin, 2019; Paul et al., 2023), enhanced learning efficiency (LeCun et al., 1989), increased interpretability (Hossain et al., 2024), accelerated learning speed (LeCun et al., 1989; Mittal et al., 2022), less interference and forgetting (Raia, 2020), forward knowledge transfer (Andle et al., 2023; Raia, 2020; Yildirim et al., 2023), and compositionality (Pfeiffer et al., 2024).
@@
Early work on sparsity in neural networks focused on simple methods such as Dropout (Srivastava et al., 2014) and L1 regularization (Tibshirani, 1996; Ng, 2004). Recent research has explored more sophisticated approaches to selectively utilize neural network parameters (Shazeer et al., 2017; Bengio, 2013). These methods generally utilize a routing or gating function (Rosenbaum et al., 2017; 2019; Shazeer et al., 2017; Pfeiffer et al., 2024), which decides which parameters or sub-networks of the model to activate based on the input. The resulting input-dependent sub-networks have been referred to as “experts” in the context of Mixture of Experts (MoE) architectures (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017), which typically consist of multiple individual parameters grouped together (e.g., a whole MLP or a transformer). Many works also explore sparsity at a much finer level of granularity, including individual neurons (Xu et al., 2024), CNN filters (Chen et al., 2019), or even single parameters (Mallya et al., 2018; Mallya & Lazebnik, 2018).
@@
However, existing sparse methods have limitations, which we would summarize as five key concerns: Firstly, most approaches rely on trainable gating functions (Shazeer et al., 2017; Mostafa & Wang, 2019; Rasmussen & Ghahramani, 2001; Li et al., 2023; Rahaman et al., 2021; Chen et al., 2019; Shazeer et al., 2018; Zhou et al., 2022; Gururangan et al., 2022; Lin et al., 2021; Ba & Frey, 2013; Bengio et al., 2016; Fedus et al., 2022; Mallya et al., 2018; Mallya & Lazebnik, 2018; Fernando et al., 2017; Keshari et al., 2018). This design choice is problematic for several reasons, including forgetfulness in continual learning (Pfeiffer et al., 2024; Raia, 2020), representation collapse (i.e., degenerate experts; Chen et al., 2023; Pfeiffer et al., 2024), complex training procedures (Rosenbaum et al., 2019), and other issues (Rosenbaum et al., 2017; Pfeiffer et al., 2024; Rosenbaum et al., 2019). Moreover, using non-trainable routing functions can be more effective (Mittal et al., 2022; Muqeeth et al., 2022). Secondly, many state-of-the-art systems employ architectures based on non-overlapping experts, where the experts do not share parameters (Shazeer et al., 2017; Pfeiffer et al., 2024). This design choice limits the transfer of previous knowledge to future tasks, also known as forward transfer, and can lead to redundancies; overlap can also be beneficial (French, 1993; Maini et al., 2023). Thirdly, even when experts overlap, it is unclear whether models can effectively learn to map similar inputs to the same experts, potentially resulting in redundancies (Chen et al., 2023) or interference (Pfeiffer et al., 2024). Furthermore, the routing mechanism often lacks interpretability, making it difficult to understand why specific inputs are routed to specific experts (Pfeiffer et al., 2024). Fourthly, many existing methods require input or task IDs to determine which mask to apply (Mallya et al., 2018; Yang et al., 2020; Masse et al., 2018; Maini et al., 2023; Pes et al., 2024; Mittal et al., 2022; Muqeeth et al., 2022; Kang et al., 2024; Wortsman et al., 2020), which can be restrictive, as meta information about inputs is rarely available in real-world applications (Aljundi et al., 2019; Ye & Bors, 2022; Wang et al., 2022). Lastly, the number of experts in current systems is limited, often ranging from a few to a couple of thousand (Shazeer et al., 2017; Jiang et al., 2024), which may not be sufficient for complex tasks (Rasmussen & Ghahramani, 2001).
@@
In this paper we introduce Conditionally Overlapping Mixture of ExperTs (COMET), a general deep learning method that induces a modular, sparse architecture in neural networks, with a number of important properties. First, COMET uses a non-trainable gating function, eliminating the need for iterative pruning or continuous sparsification. Instead, we employ a $k$-winner-take-all cap operation, inspired by the brain’s efficient use of a limited number of active cells via lateral inhibition. This design choice is both biologically motivated and automatically determined through fixed random projections. Because of it, COMET does not require fixed specialization of each network module, or advanced knowledge of the combination of modules required for each task, enabling more flexibility and adaptability. Second, the number of possible experts in COMET is exponential in the model size, exceeding the limit of a few thousand in recent work, to effectively tackle more complex tasks. Third, these experts overlap based on unsupervised information from input similarities. This, in turn, enables positive knowledge transfer between items, resulting in faster learning and improved generalization. It does this without increasing the number of trainable parameters, or requiring input or task IDs to determine which mask to apply.
@@
We validate our approach through experiments on seven diverse tasks, including image classification, language modeling, and regression, demonstrating that our method is applicable to many popular model architectures such as vision transformers, MLP-mixers, GPTs, and standard MLPs, and consistently provides improved performance.
@@ 2 RELATED WORK
Sparsity Sparsity in deep learning refers to systems where not all parameters are active or participating in the computational graph. Fixed sparsity, such as L1 regularization, intends to minimize the number of active parameters through optimizing a continuous loss function. Variable sparsity includes various methods, such as randomness as in Dropout (Srivastava et al., 2014), and input-dependent, such as in MoE (Shazeer et al., 2017). Notably, continuously optimizing deep networks to be sparse is an NP-hard problem (Jacobs & Burkholz, 2024), and pre-defined sparse architectures, such as MoE, can be restrictive. To circumvent that, we propose sparsification using a biologically motivated approach of random projection followed by a cap operation, which activates the strongest cells in the network, similar to the sensory system of fruit flies (Bruhin & Davies, 2022).
@@
Modularity Modularity is related to sparsity, where information is conditionally routed to a subset of the network’s parameters (Pfeiffer et al., 2024; Rosenbaum et al., 2017). Modular methods can be split into two. Those with trainable routing functions, including the standard MoE (Shazeer et al., 2017), adaptively determine the active parameters during training. Although this approach is widely used, it can lead to representation collapse, forgetfulness, and redundant computation (Pfeiffer et al., 2024). Secondly, fixed routing function methods, where the decision on active parameters is fixed throughout training, have been shown to be more effective (Mittal et al., 2022; Muqeeth et al., 2022). However, these methods typically require prior knowledge of module specialization (Pfeiffer et al., 2024; Muqeeth et al., 2022; Mittal et al., 2022), which is often not available in practice. In contrast, our work demonstrates that module specialization can be achieved in a fully unsupervised manner using fixed random projections. Unlike previous studies that employed fixed random projections as a routing function, either with non-overlapping experts (Roller et al., 2021; Chen et al., 2023) or with reduced performance (Bruhin & Davies, 2022), our approach focuses on the more challenging setting of overlapping experts and achieves large performance gains across diverse tasks. Additionally, our expert overlap correlates with input similarity, making it more likely that parameters are shared between similar inputs. This, in turn, enables positive knowledge transfer between items, resulting in faster learning and improved generalization. Moreover, as the experts overlap, our approach benefits from input-dependent sparsity without an explicitly modular architecture (like in MoE).
@@
Conditional Computation Conditional computation integrates sparsity and modularity, in that parameters are dropped in a learned and optimized manner, rather than randomly and independently (Bengio, 2013). A notable example is the standard MoE framework (Shazeer et al., 2017). Our approach diverges from previous work by uniquely integrating several desirable properties, including fixed routing, overlapping expert assignments based on input similarities, mask determination without requiring meta information, and scalability to an exponentially large number of experts.
@@
Dynamic Neural Networks Dynamic neural networks (Han et al., 2021) share similarities with conditional computation. Unlike traditional MoE approaches, dynamic neural networks do not have explicitly defined experts. Instead, they dynamically select units, layers, or components from the main model for each input (Han et al., 2021). Our approach has parallels with some dynamic neural network methods, such as Piggyback (Mallya et al., 2018) and PackNet (Mallya & Lazebnik, 2018), which also do not define explicit experts. However, our method differs in two key ways. Firstly, it does not require meta information to determine the expert mask. Secondly, it ensures that any given input always maps to the same expert throughout both training and inference. This contrasts with architectures that use trainable gating functions, where the same input can activate different experts or representations at different stages of learning, which can be seen as a “moving target”.
@@ 3 CONDITIONALLY OVERLAPPING MIXTURE OF EXPERTS (COMET)
COMET’s routing network is a second MLP with the same shape, defined by random weight matrices ${\bm{V}}_{\ell}$ (for simplicity we omit bias parameters). We sample ${\bm{V}}_{\ell}$ from the same distribution used for initializing ${\bm{W}}_{\ell}$ ($U(-N_{\ell-1}^{-1/2},N_{\ell-1}^{-1/2})$ in our experiments). We denote this network’s pre-activations and activations as ${\bm{c}}_{\ell}$ and ${\bm{z}}_{\ell}$ (analogous to ${\bm{a}}_{\ell}$ and ${\bm{x}}_{\ell}$ in the backbone network), with input ${\bm{z}}_{0}={\bm{x}}_{0}$. The computation of the routing network is similar to that of the backbone MLP, except that at each layer it computes a binary vector ${\bm{m}}_{\ell}$ that is then used to mask the activations in both networks. The mask is computed using a capping function $C_{k}$ that identifies the top $k$ elements of a vector:
@@
COMET layers differ from standard MoE (Shazeer et al., 2017) layers in two major ways:
@@
    Architecturally: Whereas a layer in a standard layered MoE architecture consists of $e$ experts and a gating network, a COMET layer contains a random, non-trainable matrix and a $k$-winner-take-all cap operation. Note that a COMET layer does not contain pre-defined experts. Instead, it dynamically modifies the computation of the MLP to activate only a subset of its parameters; this subset can be seen as an implicit expert. This input-dependent masking at the neuron level of each layer results in a maximum number of experts that is exponential in the model size, specifically (${N\choose k}$. For instance, a small layer with 1000 neurons and a moderate sparsity level of $10\%$ yields an upper bound of (${1000\choose 100}\approx 6.38e^{139}$ experts. Therefore in practical settings every input will have its own expert.
@@
    In the way the information is passed: Whereas in MoE the input to the gating network is the same as the input to each of the MLP layers, in a COMET architecture the inputs are distinct (except for the first layer of the network). That is, the routing network operates independently of the backbone network. This ensures that the same inputs always map to the same implicit expert throughout both training and inference.

### v4

@@ Abstract
The evolution of biological neural systems has led to both modularity and sparse coding, which enables energy efficiency and robustness across the diversity of tasks in the lifespan. In contrast, standard neural networks rely on dense, non-specialized architectures, where all model parameters are simultaneously updated to learn multiple tasks, leading to interference. Current sparse neural network approaches aim to alleviate this issue but are hindered by limitations such as 1) trainable gating functions that cause representation collapse, 2) disjoint experts that result in redundant computation and slow learning, and 3) reliance on explicit input or task IDs that limit flexibility and scalability. In this paper we propose Conditionally Overlapping Mixture of ExperTs (COMET), a general deep learning method that addresses these challenges by inducing a modular, sparse architecture with an exponential number of overlapping experts. COMET replaces the trainable gating function used in Sparse Mixture of Experts with a fixed, biologically inspired random projection applied to individual input representations. This design causes the degree of expert overlap to depend on input similarity, so that similar inputs tend to share more parameters. This results in faster learning per update step and improved out-of-sample generalization. We demonstrate the effectiveness of COMET on a range of tasks, including image classification, language modeling, and regression, using several popular deep learning architectures. Code can be found here: https://github.com/Shaier/COMET.git.
@@ 1 Introduction
In recent years, there has been a trend towards developing increasingly larger models (OpenAI, 2023a; b; Fedus et al., 2022; Shuster et al., 2022; Chowdhery et al., 2022), driven by the understanding that a neural network’s learning capacity depends on its number of parameters (Shazeer et al., 2017). This approach has yielded impressive results in various fields, including computer vision (Dosovitskiy et al., 2021; Kirillov et al., 2023) and language modeling (OpenAI, 2023b; Chowdhery et al., 2022). However, with such large size come difficulties, including increased training costs and growing requirements for large amounts of memory and storage.
@@
One approach to mitigating some of these challenges is sparsity, where a subset of the model’s parameters is selectively utilized in the computational graph. This concept of sparsity has been widely explored in machine learning (Jacobs et al., 1991; Jordan & Jacobs, 1993; LeCun et al., 1989; Zhou et al., 2019; Hoefler et al., 2021; Shazeer et al., 2017; Bair et al., 2024). Researchers have observed significant benefits of sparsity, including reduced inference costs (Han et al., 2016; Shazeer et al., 2017), improved generalization capabilities (LeCun et al., 1989; Jacobs & Burkholz, 2024; Frankle & Carbin, 2019; Paul et al., 2023), enhanced learning efficiency (LeCun et al., 1989), accelerated learning speed (LeCun et al., 1989; Mittal et al., 2022), less interference and forgetting (Raia, 2020), forward knowledge transfer (Andle et al., 2023; Raia, 2020; Yildirim et al., 2023), and compositionality (Pfeiffer et al., 2024).
@@
Early work on sparsity in neural networks focused on simple methods such as Dropout (Srivastava et al., 2014) and L1 regularization (Tibshirani, 1996; Ng, 2004). These and subsequent works explore sparsity at a fine level of granularity, including single parameters (Mallya et al., 2018; Mallya & Lazebnik, 2018), individual neurons (Xu et al., 2024), or CNN filters (Chen et al., 2019). Other research has explored sparsity at the level of whole networks or sub-networks within the mixture of experts (MoE) framework (Shazeer et al., 2017; Bengio, 2013). These methods generally utilize a routing or gating function (Rosenbaum et al., 2017; 2019; Shazeer et al., 2017; Pfeiffer et al., 2024), which decides which parameters or sub-networks of the model to activate based on the input.
@@
However, existing sparse methods have limitations, which we would summarize as five key concerns: Firstly, most approaches rely on trainable gating functions (Shazeer et al., 2017; Mostafa & Wang, 2019; Rasmussen & Ghahramani, 2001; Li et al., 2023; Rahaman et al., 2021; Chen et al., 2019; Shazeer et al., 2018; Zhou et al., 2022; Gururangan et al., 2022; Lin et al., 2021; Ba & Frey, 2013; Bengio et al., 2016; Fedus et al., 2022; Mallya et al., 2018; Mallya & Lazebnik, 2018; Fernando et al., 2017; Keshari et al., 2018). This design choice is problematic for several reasons, including forgetfulness in continual learning (Pfeiffer et al., 2024; Raia, 2020), representation collapse (i.e., degenerate experts; Chen et al., 2023; Pfeiffer et al., 2024), complex training procedures (Rosenbaum et al., 2019), and other issues (Rosenbaum et al., 2017; Pfeiffer et al., 2024; Rosenbaum et al., 2019). Moreover, using non-trainable routing functions can be more effective (Mittal et al., 2022; Muqeeth et al., 2022). Secondly, many state-of-the-art systems employ architectures based on disjoint experts that do not share parameters (Shazeer et al., 2017; Pfeiffer et al., 2024). This design choice can lead to redundancies and may limit generalization; overlap can also be beneficial (French, 1993; Maini et al., 2023). Thirdly, even when experts overlap, it is unclear whether models can effectively learn to map similar inputs to the same experts, potentially resulting in redundancies (Chen et al., 2023) or interference (Pfeiffer et al., 2024). Fourthly, many existing methods require input or task IDs to determine which mask to apply (Mallya et al., 2018; Yang et al., 2020; Masse et al., 2018; Maini et al., 2023; Pes et al., 2024; Mittal et al., 2022; Muqeeth et al., 2022; Kang et al., 2024; Wortsman et al., 2020), which can be restrictive, as meta-information about inputs is rarely available in real-world applications (Aljundi et al., 2019; Ye & Bors, 2022; Wang et al., 2022). Lastly, the number of experts in current systems is limited, often ranging from a few to a couple of thousand (Shazeer et al., 2017; Jiang et al., 2024), which may not be sufficient for complex tasks (Rasmussen & Ghahramani, 2001).
@@
In this paper we introduce Conditionally Overlapping Mixture of ExperTs (COMET), a general deep learning method that induces a modular, sparse architecture in neural networks, with a number of important properties. First, COMET uses a non-trainable gating function, eliminating the need for iterative pruning or continuous sparsification. Instead, we employ a fixed random projection followed by a $k$-winner-take-all cap operation, inspired by the brain’s efficient use of a limited number of active cells via lateral inhibition. As in the brain, these mechanisms combine to produce sparse representations with overlap that depends on input similarity (Bruhin & Davies, 2022). Second, COMET does not require fixed specialization of sub-networks or advance knowledge of the active neurons required for each task, enabling more flexibility and adaptability. Third, the number of possible experts in COMET is exponential in the model size, exceeding the limit of a few thousand in recent work, to effectively tackle more complex tasks. Fourth, these experts overlap based on unsupervised information from input similarities. This yields faster learning and improved generalization. It does this without increasing the number of trainable parameters, or requiring input or task IDs to determine which mask to apply.
COMET integrates concepts from diverse research areas into a concise framework: fixed random projection and $k$-winner-take-all from neuroscience, routing functions from modular neural networks, expert-based approaches from the MoE literature, the notion of implicit experts from dynamic neural networks, the integration of sparsity and modularity from conditional computation, input-dependent masking from various deep learning areas, and the importance of active parameter overlap from continual learning. The present paper focuses on learning and out-of-sample generalization in single tasks, but we conjecture COMET’s input-dependent sparsity will also yield advantages for settings involving multiple tasks, including transfer learning, continual learning, and robustness to catastrophic forgetting. These will be tested in future work, although we report a preliminary test of transfer learning in section A.10.
@@
We validate our approach through experiments on seven diverse tasks, including image classification, language modeling, and regression, demonstrating that our method is applicable to many popular model architectures such as vision transformers, MLP-mixers, GPTs, and standard MLPs, and consistently provides improved performance. Our code will be linked here upon paper acceptance.
@@ 2 Related Work
Sparsity Sparsity in deep learning refers to systems where not all parameters are active or participating in the computational graph. Fixed sparsity, such as L1 regularization, intends to minimize the number of active parameters through optimizing a continuous loss function. Variable sparsity includes various methods, such as randomness as in Dropout (Srivastava et al., 2014), and input-dependent, such as in sparse MoE (Shazeer et al., 2017). Notably, continuously optimizing deep networks to be sparse is an NP-hard problem (Jacobs & Burkholz, 2024), and pre-defined sparse architectures can be restrictive. To circumvent that, we propose sparsification using a biologically motivated approach of random projection followed by a cap operation, which activates the strongest cells in the network, similar to the sensory system of fruit flies (Bruhin & Davies, 2022).
@@
Modularity Modularity is related to sparsity, where information is conditionally routed to a subset of the network’s parameters (Pfeiffer et al., 2024; Rosenbaum et al., 2017). Modular methods can be split into two types. Firstly, those with trainable routing functions, including the sparse MoE (Shazeer et al., 2017), adaptively determine the active parameters during training. Although widely used, this approach can lead to representation collapse, forgetfulness, and redundant computation (Pfeiffer et al., 2024). Secondly, fixed routing function methods have been shown to be more effective (Mittal et al., 2022; Muqeeth et al., 2022). However, these methods typically require prior knowledge of module specialization (Pfeiffer et al., 2024; Muqeeth et al., 2022; Mittal et al., 2022), which is often not available in practice. In contrast, our work demonstrates that module specialization can be achieved in a fully unsupervised manner using fixed random projections. Unlike previous studies that employed fixed random projections as routing functions, either with non-overlapping experts (Roller et al., 2021; Chen et al., 2023) or with reduced performance (Bruhin & Davies, 2022), our approach focuses on the more challenging setting of overlapping experts and achieves large performance gains across diverse tasks. Additionally, our expert overlap correlates with input similarity, making it more likely that parameters are shared between similar inputs. This, in turn, enables positive knowledge transfer between items, resulting in faster learning and improved generalization. Moreover, as the experts overlap, our approach benefits from input-dependent sparsity without an explicitly modular architecture (like in MoE).
@@
Conditional Computation Conditional computation integrates sparsity and modularity, in that parameters are dropped in a learned and optimized manner, rather than randomly and independently (Bengio, 2013). A notable example is the sparse MoE framework (Shazeer et al., 2017). Our approach diverges from previous work by uniquely integrating several desirable properties, including fixed routing, overlapping expert assignments based on input similarities, mask determination without requiring meta-information, and scalability to an exponentially large number of experts.
@@
Dynamic Neural Networks Dynamic neural networks (Han et al., 2021) share similarities with conditional computation. Unlike traditional MoE approaches, dynamic neural networks do not have explicitly defined experts. Instead, they dynamically select units, layers, or components from the main model for each input (Han et al., 2021). Our approach has parallels with some dynamic neural network methods, such as Piggyback (Mallya et al., 2018) and PackNet (Mallya & Lazebnik, 2018), which also do not define explicit experts. However, our method differs in two key ways. Firstly, it does not require meta-information to determine the expert mask. Secondly, it ensures that any given input always maps to the same expert throughout both training and inference. This contrasts with architectures that use trainable gating functions, where the same input can activate different experts or representations at different stages of learning, which can be seen as a “moving target”.
The example-tied dropout scheme of Maini et al. (2023) is another dynamic neural network in which neurons are partitioned into a “generalization” pool that is active for all inputs and a “memorization” pool for which a small subset is active for each input. The assignment of memorization neurons to each input example is random and fixed throughout training. Our approach differs from example-tied dropout in that similar inputs share more active neurons, which facilitates generalization.
The model superposition of cheung2019superposition applies a task-dependent orthogonal transformation to the activation vector at each layer. Assuming the input distribution has effective dimensionality less than that of the activation space (i.e., the layer width), these transformations map different tasks into different subspaces allowing them to be learned semi-independently. Our input-dependent masking also restricts the representation of each example to a subspace, but it does not require task IDs and it leads similar inputs to have more overlapping representations.
3 Mixture of Experts and Input-dependent Masking
A standard MoE architecture involves a disjoint set of experts and a gate that combines their predictions (Jacobs et al., 1991). For example in the sparse MoE framework proposed by Shazeer et al. (2017), each MoE module consists of $n$ expert networks, $E_{1},...,E_{n}$, and a gating network, $G$, that outputs a sparse $n$-dimensional vector of mixture weights. The gating and expert networks are all trainable, each with its own set of parameters. The prediction for an input ${\bm{x}}$ is $\sum_{i=1}^{n}G({\bm{x}})_{i}E_{i}({\bm{x}})$.
Separately, several recent works have proposed versions of input-dependent masking (Maini et al., 2023; Mallya et al., 2018; Mallya & Lazebnik, 2018; Yang et al., 2020; hung2019compactingpickinggrowingunforgetting). The general framework involves a network with $n$ neurons and a masking function $m\!:\mathcal{X}\to\{0,1\}^{n}$ (where $\mathcal{X}$ is the input space). In processing an example ${\bm{x}}$, the network’s activations are multiplied elementwise with $m({\bm{x}})$. Thus the prediction for ${\bm{x}}$ is $F_{m({\bm{x}})}({\bm{x}})$ where $F_{m({\bm{x}})}$ is the function computed by the sub-network corresponding to the mask $m({\bm{x}})$.
Combining these two lines of work, we propose to view the sub-networks defined by input-dependent masking as overlapping experts. Any two experts will typically share many active neurons, and hence weights. This is in contrast to standard MoE where the experts learn disjoint sets of parameters. In the overlapping MoE framework, every subset of the full network is a (potential) expert. The sub-network $F_{m({\bm{x}})}$ is an expert for ${\bm{x}}$, and it is also a partial expert for any other ${\bm{x}}^{\prime}$ to the degree that $m({\bm{x}})$ and $m({\bm{x}}^{\prime})$ overlap, as determined by the inner product $m({\bm{x}})^{\top}m({\bm{x}}^{\prime})$. In the overlapping MoE framework, the gating network $G$ is replaced by the masking function $m$.
We further propose that similar inputs should map to similar (i.e., more overlapping) experts. This will facilitate generalization because what is learned about one input will selectively generalize to similar inputs. The next section explains how COMET achieves this property using biologically inspired fixed random projections (${\bm{V}}_{\ell}$ in eq. 6) and $k$-winner-take-all capping ($C_{k_{\ell}}$ in eq. 7).
@@ 4 Conditionally Overlapping Mixture of Experts (COMET)
COMET’s routing network is a second MLP with the same shape, defined by random weight matrices ${\bm{V}}_{\ell}$ (for simplicity we omit bias parameters). We sample ${\bm{V}}_{\ell}$ from the same distribution used for initializing ${\bm{W}}_{\ell}$ ($U(-N_{\ell-1}^{-1/2},N_{\ell-1}^{-1/2})$ in our experiments). We denote this network’s pre-activations and activations as ${\bm{c}}_{\ell}$ and ${\bm{z}}_{\ell}$ (analogous to ${\bm{a}}_{\ell}$ and ${\bm{x}}_{\ell}$ in the backbone network), with input ${\bm{z}}_{0}={\bm{x}}_{0}$. The computation of the routing network is similar to that of the backbone MLP, except that at each layer it computes a binary vector ${\bm{m}}_{\ell}$ that is then used to mask the activations in both networks. The mask is computed using a $k$-winner-take-all capping function $C_{k}$:
@@
This input-dependent masking results in a maximum number of experts that is exponential in the model size at each layer, specifically (${N_{\ell}\choose k}$. Therefore in practical settings every input will have its own expert. One consideration for this calculation might be interference among experts. Previous work has studied how multiple models can be superposed within one network (cheung2019superposition), and elhage2022toy show a layer with $N_{\ell}$ neurons can hold $O(e^{N})$ representations that are pairwise orthogonal within a certain finite tolerance, thus minimizing interference. However, more important for the present work is that overlap between models is a desired property because it promotes generalization between similar inputs.
COMET layers differ from sparse MoE (Shazeer et al., 2017) layers in two major ways:
@@
    Architecturally: Whereas a layer in a standard layered MoE architecture consists of $n$ experts and a gating network, a COMET layer contains a random, non-trainable matrix and a $k$-winner-take-all cap operation. Instead of pre-defined experts, COMET layer modifies the computation of the MLP to activate only a subset of its parameters contingent on the input; this subset can be seen as an implicit expert.
@@
    In the way the information is passed: Sparse MoE is applied in layers with a new gating network at each layer, which takes as input the backbone activation at the previous layer. Thus the gating and backbone networks at each layer take the same input. In a COMET architecture the routing network operates independently of the backbone network, so the inputs to the two are distinct (except for the first layer of the network). This ensures that a given example maps to the same implicit expert throughout both training and inference.
Several other important differences between existing approaches and COMET are: 1) COMET’s gating network does not require training; 2) COMET does not require fixed specialization of each network module, or advanced knowledge of the combination of modules required for each task; 3) the experts in COMET overlap based on unsupervised information from input similarities; 4) COMET does not require input or task IDs to determine which mask to apply; 5) the number of possible experts in COMET is exponential in the model size.